If 2x-3(x+4)= -5, then x =
Answer by algebrahouse.com(1659)

"If 2x-3(x+4)= -5, then x ="

2x - 3(x + 4)= -5
2x - 3x - 12 = -5 {used distributive property}
-x - 12 = -5 {combined like terms}
-x = 7 {added 12 to both sides}
x = -7 {divided both sides by -1}
www.algebrahouse.com